# LIR: A Lightweight Baseline for Image Restoration

### v2

@@ 5.1 Conclusion
In this paper, we introduce a Lightweight Image Restoration network called LIR, which aims to efficiently remove degradation. Firstly, the problem that local and global residual connections propagate the degradation throughout the entire network is addressed through the use of the Transpose operation and a simple structural design. In addition, we propose the Adaptive Filter (AF) to adaptively extract high-frequency information, enhance object contours, and eliminate degradation in various IR tasks. Our AF achieves promising results while requiring minimal parameters and computations, as demonstrated in our ablation studies (Tab. 5). To capture global information while keeping small parameters and low computational requirements, we introduce the Patch Attention (PA) module by leveraging an MLP and patch embedding strategy.
5.2 Limitation
@@
For the deraining task, LIR achieves comparable performance with the state-of-the-art method on PSNR and the best performance on SSIM with smaller parameter numbers and computations; For denoising and dehazing tasks, LIR also achieves comparable performance to the state-of-art method. In addition, LIR achieves state-of-the-art visual results on deraining and denoising tasks. However, as Tab. 4 shows, LIR is less prominent on deblurring tasks than deraining and denoising. We believe that this is because, in the deblurring task, object outlines tend to be stretched out very long. While our Adaptive Filter is very effective when dealing with static object contours, it has limitations when dealing with dynamic contours caused by fast-moving objects.

### v3

@@ 5.1. Conclusion
In this paper, we introduce a Lightweight Image Restoration network called LIR, which aims to efficiently remove degradation. Firstly, the problem that local and global residual connections propagate the degradation throughout the entire network is addressed through the use of the Transpose operation and an ingenious structural design. In addition, we propose the Adaptive Filter to adaptively extract high-frequency information, enhance object contours, and eliminate degradation in various IR tasks. Our Adaptive Filter achieves promising results while requiring minimal parameters and computations, as demonstrated in our ablation studies (Section 5). To capture global information while keeping small parameters and low computational requirements, we introduce the Patch Attention module by leveraging an MLP and patch embedding strategy. On the deraining task, our LIR achieves the state-of-the-art Structure Similarity Index Measure and comparable performance to state-of-the-art models on Peak Signal-to-Noise Ratio. For denoising, dehazing, and deblurring tasks, LIR also achieves a comparable performance to state-of-the-art models with only about 30% size.
5.2. Limitation
@@
LIR achieves good performance on deraining, denoising and dehazing tasks and better visual results. However, as Figure 4 shows, LIR is less prominent on deblurring tasks than deraining and denoising. We believe that this is because, in the deblurring task, object outlines tend to be stretched out very long. While our Adaptive Filter is very effective when dealing with static object contours, it has limitations when dealing with dynamic contours caused by fast-moving objects.